OpenStudy (anonymous):

A construction cone has a diameter of 8 inches and a height of 16 inches. What is the volume of the construction cone? Round your answer to the nearest hundredth.

OpenStudy (lgbasallote):

volume of cone = 1/3 pi r^2 h since you have diameter turn it into radius first radius = diameter/2 radius = 8/2 radius = 4 now put this into the equation \[V = \frac 13 \pi r^2 h\] \[V = \frac 13 \pi \times (4)^2 \times 16\] got it?